# LoBAM: LoRA-Based Backdoor Attack on Model Merging

### v2

@@ Abstract
Model merging is an emerging technique that integrates multiple models fine-tuned on different tasks to create a versatile model that excels in multiple domains. This scheme, in the meantime, may open up backdoor attack opportunities where one single malicious model can jeopardize the integrity of the merged model. Existing works try to demonstrate the risk of such attacks by assuming substantial computational resources, focusing on cases where the attacker can fully fine-tune the pre-trained model. Such an assumption, however, may not be feasible given the increasing size of machine learning models. In practice where resources are limited and the attacker can only employ techniques like Low-Rank Adaptation (LoRA) to produce the malicious model, it remains unclear whether the attack can still work and pose threats. In this work, we first identify that the attack efficacy is significantly diminished when using LoRA for fine-tuning. Then, we propose LoBAM, a method that yields high attack success rate with minimal training resources. The key idea of LoBAM is to amplify the malicious weights in an intelligent way that effectively enhances the attack efficacy. We demonstrate that our design can lead to improved attack success rate through both theoretical proof and extensive empirical experiments across various model merging scenarios. Moreover, we show that our method has strong stealthiness and is difficult to detect.
@@ 1 Introduction
In this paper, we address this gap by introducing a novel attack algorithm, LoBAM, which to our knowledge is the first method that effectively exposes the security risks of the backdoor attack against model merging in low-resource scenarios. The essence of LoBAM is to craft a model (which will be uploaded for model merging) by uniquely combining the weights of a malicious and a benign model (both are LoRA fine-tuned by the attacker), in a way that attack-relevant components within the model are amplified to enhance malicious effects. Our design is inspired by certain findings about LoRA [26] and is further backed up by our mathematical proof which guarantees increased attack success rate when applying the proposed LoBAM.
We conduct extensive experiments to validate our method. Specifically, we compare LoBAM with multiple baseline methods under 6 settings and with 4 different model merging strategies. Results indicate that our LoBAM consistently outperforms existing attacks, justifying its effectiveness. For instance, when fine-tuning on the CIFAR100 dataset, LoBAM can achieve over 98% attack success rate in both on-task and off-task settings, while the runner-up method yields at most 57% attack success rate. We also examine whether LoBAM could be detected during the model merging process. To this end, we perform a t-SNE analysis [40, 2], which is commonly used for low-dimensional visualization and detection of malicious models. The results reveal that the latent space distributions of benign and malicious models are nearly indistinguishable, demonstrating that our proposed attack remains stealthy.
@@
  • We propose a novel and computationally efficient attack method against model merging, which is supported by rigorous mathematical proof.
  • With extensive experiments, we demonstrate that the proposed method delivers outstanding attack performance across diverse scenarios while maintaining a high level of stealth against detection.
@@ 2.3 Backdoor Attacks on Model Merging
Backdoor attacks [12, 35, 51] aim to manipulate the training process of machine learning models so that the final model exhibits specific, targeted misbehavior when the input is attached with a particular trigger. While most works studying backdoor attacks focus on centralized or single-model settings [12, 35], BadMerging [51] designs a backdoor attack that targets model merging, where the final merged model can be compromised with the malicious model uploaded by the attacker. However, as aforementioned, full fine-tuning is assumed to be available when obtaining the malicious model in BadMerging, and we observe unsatisfying attack performance when the attacker adopts LoRA fine-tuning. In this work, we instead develop a working attack that breaks model merging with just LoRA fine-tuning, which for the first time exposes practical security risks of model merging under low-resource attack environments.
@@ 3.3 Attacker’s Knowledge
In our attack scenario, the attacker has no prior knowledge of the training data used by benign users, the benign models to be merged, or the merging algorithm. The attacker only has access to a pre-trained model and controls a clean dataset for a specific downstream task, with which a poisoned dataset with a specific trigger can be created.
If the attacker aims to execute an off-task attack, they also possess a few images of the targeted class in addition to the aforementioned datasets [51]. For instance, if the attacker employs CIFAR100 datasets for fine-tuning, and their objective is to cause the merged model to misclassify images as ‘stop’ when seeing a trigger-attached image from GTSRB, the attacker would only need a few images labeled as ‘stop,’ without requiring any other images from GTSRB.
@@ 4 Our Attack
In this section, we first motivate our method by showing and analyzing that LoRA fine-tuning degrades the attack performance of the merging attack. Then, we present and discuss the proposed method in detail. Lastly, we theoretically prove that our method can lead to improved attack success rate.
@@ 4.2 LoBAM
4.3 Mathematical Analysis and Proof
In this section, we theoretically prove that our proposed LoBAM can achieve an improved attack success rate. Let $\theta_{\text{pre}}$ represent the pre-trained model and suppose the $k$-th user is the attacker. $\Delta\theta_{k_{m}}^{\prime}$ and $\Delta\theta_{k_{b}}^{\prime}$ represent the weight updates of the malicious model ($\theta_{\text{malicious}}$) and the benign model ($\theta_{\text{benign}}$) fine-tuned by the attacker, respectively. $\Delta\theta_{i}$ ($i\neq k$) stands for the weight updates of the $i$-th benign model fine-tuned by each of the other benign users. $N$ is the total number of models to merge. $g(\cdot)$ represents the attack success rate upon a certain model.
Before introducing the theoretical analysis, we first present our technical assumption on convexity, which is widely adopted in existing literature [11, 10, 15].
Assumption 1.
The function $g$ is $\mu$-strongly convex, i.e., for all $\theta_{1},\theta_{2}\in\Theta$, one has
$g(\theta_{2})\geq g(\theta_{1})+\left\langle{\nabla g(\theta_{1}),\theta_{2}-\theta_{1}}\right\rangle+\frac{\mu}{2}{\left\|{\theta_{2}-\theta_{1}}\right\|^{2}}.$
Theorem 1.
Suppose Assumptions 1 holds. We define $Y=\theta_{\text{pre}}+\frac{1}{N}\left(\sum_{i=1,i\neq k}^{N}\Delta\theta_{i}+\Delta\theta_{k_{m}}^{\prime}\right)$, which is the merged model weights when our attack is not applied (i.e., directly uploading the fine-tuned malicious model without applying Equation 3). Choose $\mathcal{G}=\left\|\nabla g^{T}\left(Y\right)\right\|$. When $\quad\lambda>1+\frac{\mathcal{G}}{\frac{\mu}{2N}\left\|\Delta\theta_{k_{m}}^{\prime}-\Delta\theta_{k_{b}}^{\prime}\right\|}$, we have
$g(X)>g(Y),$
where
$X=\theta_{\text{pre}}+\frac{1}{N}\left(\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\Delta\theta_{i}+\lambda\left(\Delta\theta_{k_{m}}^{\prime}-\Delta\theta_{k_{b}}^{\prime}\right)+\Delta\theta_{k_{b}}^{\prime}\right)$
represents the merged model weights with our attack being applied.
We leave the proof to Appendix A. This finding directly points out that the attacker can achieve a higher attack success rate using LoBAM with $\lambda>1$ than without applying our method.
@@ C MNIST, Cars196, ImageNet100, STL10, EuroSAT, GTSRB
Parameter setting:
When the attacker constructs the malicious model, we set $r=8$ for LoRA and $\lambda_{\text{min}}=4$ and $\lambda_{\text{max}}=10$ for Algorithm 2. Later we will show the results under various $r$ and $\lambda$ for our method.
@@ 5.3 Safety Detection and Defense
To ensure that our proposed LoBAM remains undetectable while uploading the malicious model to the open platform, we perform t-SNE analysis [40, 2] on both benign and malicious models. Unlike PCA [28] or sub-sampling [31] techniques, t-SNE excels at preserving the original data distributions in a lower-dimensional space, making it a superior choice for identifying and defending against malicious activity according to previous research [52, 39, 30].
In our experiment, we analyze a total of 80 models—60 benign models fine-tuned on various tasks and 20 malicious models created using LoBAM, also derived from diverse datasets. We apply t-SNE to reduce the dimensionality of model parameters from each layer to three dimensions for visualization purposes. The visualization results for all the layers are depicted in Figure 4 in the Appendix, demonstrating that the parameters of both benign and malicious models are indistinguishable within the latent space. This finding provides compelling evidence of the robust concealment capabilities of our proposed attack, confirming its evasion from detection and defense.
@@ 6 Conclusion
In this paper, we discovered that existing backdoor attacks on model merging become ineffective due to attackers’ limited computational resources and the resulting reliance on LoRA for fine-tuning pre-trained models. Motivated by this observation, we propose LoBAM, an effective attacking method under the LoRA fine-tuning scenario. LoBAM strategically combines the weights of a malicious and a benign model—each LoRA fine-tuned by the attacker—to amplify attack-relevant components, enhancing the model’s malicious efficacy when deployed in model merging. Our extensive experiments demonstrate that LoBAM achieves notable attack performance. Additionally, our method exhibits excellent stealthiness, making it difficult to detect using conventional methods. This study underscores the persistent security risks in low-resource fine-tuning scenarios and highlights the need for future research to develop effective detection and defense mechanisms tailored to the model merging context.
Appendix A Proof of Theorem 1
Let
$A=g\left(X\right)$
$B=g\left(Y\right)$
We want to prove:
$A>B\quad\text{if}\quad\lambda\quad\text{is larger than a specific value}.$
Proof:
$A$ $=g\Bigg{(}\underbrace{\theta_{\text{pre}}+\frac{1}{N}\bigg{(}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\Delta\theta_{i}+\Delta\theta_{k_{m}}^{\prime}\bigg{)}}_{Y}$
$\qquad+\underbrace{\frac{1}{N}\left(\lambda(\Delta\theta_{k_{m}}^{\prime}-\Delta\theta_{k_{b}}^{\prime})+\Delta\theta_{k_{b}}^{\prime}-\Delta\theta_{k_{m}}^{\prime}\right)}_{Z}\Bigg{)}$
Following the strong convexity of $g$, we have
$A\geq B+\nabla g^{T}(Y)Z+\frac{\mu}{2}\|Z\|^{2}.$
So we only need to guarantee:
$\nabla g^{T}(Y)Z+\frac{\mu}{2}\|Z\|^{2}>0,$
which can be guaranteed if the minimum of it is $>0$, which is:
$-\left\|\nabla g^{T}\left(Y\right)\right\|\cdot\|Z\|+\frac{\mu}{2}\|Z\|^{2}>0,$
which is equal to:
$\left\|\nabla g^{T}\left(Y\right)\right\|\leq\frac{\mu}{2}\|Z\|.$
We denote $\left\|\nabla g^{T}\left(Y\right)\right\|$ as $G$. Because
$\frac{\mu}{2}\|Z\|$ $=\frac{\mu}{2}\left\|\frac{1}{N}\left(\lambda(\Delta\theta_{k_{m}}^{\prime}-\Delta\theta_{k_{b}}^{\prime})+\Delta\theta_{k_{b}}^{\prime}-\Delta\theta_{k_{m}}^{\prime}\right)\right\|$
$\geq\frac{\lambda\mu}{2N}\left\|\Delta\theta_{k_{m}}^{\prime}-\Delta\theta_{k_{b}}^{\prime}\right\|-\frac{\mu}{2N}\left\|\Delta\theta_{k_{m}}^{\prime}-\Delta\theta_{k_{b}}^{\prime}\right\|,$
$A>B$ can be guaranteed if we have:
$\frac{(\lambda-1)\mu}{2N}\left\|\Delta\theta_{k_{m}}^{\prime}-\Delta\theta_{k_{b}}^{\prime}\right\|>\mathcal{G}$
therefore, when $\quad\lambda>1+\frac{\mathcal{G}}{\frac{\mu}{2N}\left\|\Delta\theta_{k_{m}}^{\prime}-\Delta\theta_{k_{b}}^{\prime}\right\|}$, we have $\quad A>B$.
Figure 4: The layers of malicious models and benign models in the latent space.

### v3

@@ Abstract
Model merging is an emerging technique that integrates multiple models fine-tuned on different tasks to create a versatile model that excels in multiple domains. This scheme, in the meantime, may open up backdoor attack opportunities where one single malicious model can jeopardize the integrity of the merged model. Existing works try to demonstrate the risk of such attacks by assuming substantial computational resources, focusing on cases where the attacker can fully fine-tune the pre-trained model. Such an assumption, however, may not be feasible given the increasing size of machine learning models. In practice where resources are limited and the attacker can only employ techniques like Low-Rank Adaptation (LoRA) to produce the malicious model, it remains unclear whether the attack can still work and pose threats. In this work, we first identify that the attack efficacy is significantly diminished when using LoRA for fine-tuning. Then, we propose LoBAM, a method that yields high attack success rate with minimal training resources. The key idea of LoBAM is to amplify the malicious weights in an intelligent way that effectively enhances the attack efficacy. We demonstrate that our design can lead to improved attack success rate through extensive empirical experiments across various model merging scenarios. Moreover, we show that our method is highly stealthy and is difficult to detect and defend against.
@@ 1 Introduction
In this paper, we address this gap by introducing a novel attack algorithm, LoBAM, which to our knowledge is the first method that effectively exposes the security risks of the backdoor attack against model merging in low-resource scenarios. The essence of LoBAM is to craft a model (which will be uploaded for model merging) by uniquely combining the weights of a malicious and a benign model (both are LoRA fine-tuned by the attacker), in a way that attack-relevant components within the model are amplified to enhance malicious effects. Our design is inspired by certain findings about LoRA [24].
We conduct extensive experiments to validate our method. Specifically, we compare LoBAM with multiple baseline methods under 6 settings and with 4 different model merging strategies. Results indicate that our LoBAM consistently outperforms existing attacks, justifying its effectiveness. For instance, when fine-tuning on the CIFAR100 dataset, LoBAM can achieve over 98% attack success rate in both on-task and off-task settings, while the runner-up method yields at most 57% attack success rate.
@@
  • We propose a novel and computationally efficient attack method against model merging.
  • With extensive experiments, we demonstrate that the proposed method delivers outstanding attack performance across diverse scenarios while maintaining a high level of stealth against detection and defense.
@@ 2.3 Backdoor Attacks on Model Merging
Backdoor attacks [10, 31, 49, 48] aim to manipulate the training process of machine learning models so that the final model exhibits specific, targeted misbehavior when the input is attached with a particular trigger. While most works studying backdoor attacks focus on centralized or single-model settings [10, 31], BadMerging [49] designs a backdoor attack that targets model merging, where the final merged model can be compromised with the malicious model uploaded by the attacker. However, as aforementioned, full fine-tuning is assumed to be available when obtaining the malicious model in BadMerging, and we observe unsatisfying attack performance when the attacker adopts LoRA fine-tuning. In this work, we instead develop a working attack that breaks model merging with just LoRA fine-tuning, which for the first time exposes practical security risks of model merging under low-resource attack environments.
@@ 3.3 Attacker’s Knowledge
In our attack scenario, the attacker has no prior knowledge of the training data used by benign users, the benign models to be merged, or the merging algorithm. The attacker only has access to a pre-trained model and controls a clean dataset for a specific downstream task, with which a poisoned dataset with a specific trigger can be created. If the attacker aims to execute an off-task attack, they also possess a few images of the targeted class in addition to the aforementioned datasets [49]. For instance, if the attacker employs CIFAR100 datasets for fine-tuning, and their objective is to cause the merged model to misclassify images as ‘stop’ when seeing a trigger-attached image from GTSRB, the attacker would only need a few images labeled as ‘stop,’ without requiring any other images from GTSRB.
@@ 4 Our Attack
In this section, we first motivate our method by showing and analyzing that LoRA fine-tuning degrades the attack performance of the merging attack. Then, we present and discuss the proposed method in detail.
@@ C MNIST, Cars196, ImageNet100, STL10, EuroSAT, GTSRB
Parameter setting:  When the attacker constructs the malicious model, we set $r=8$ for LoRA and $\lambda_{\text{min}}=4$ and $\lambda_{\text{max}}=10$ for Algorithm 2. Later we will show the results under various $r$ and $\lambda$ for our method.
@@ 5.2 Experimental Results
Study on the impact of different models:  In the default setting, we primarily focus on scenarios where all users use the ViT-L/14 model due to the reasons mentioned above. However, there are cases where users employ smaller models, such as ViT-B/32 or ViT-B/16, in the model merging system. To evaluate the effectiveness of LoBAM in such scenarios, we conducted experiments accordingly. As demonstrated in Table 11 and Table 12, the results reveal that LoBAM maintains excellent performance across varying models.
@@ A B C
Table 11: Attack success rate (%) on on-task and off-task scenarios on dataset combination A, B, and C when all users use ViT-B/32 in the model merging system. SA, TA, Ties, and AM are four different algorithms.
A B C
on-task off-task on-task off-task on-task off-task
SA 98.26 99.71 98.45 99.85 100 97.32
TA 100 99.93 100 100 99.31 99.82
Ties 98.05 99.57 97.52 99.94 98.16 100
AM 99.84 98.47 100 98.77 99.55 99.42
@@ 5.3 Safety Detection and Defense
Safety Detection:
To ensure that the malicious model generated by LoBAM remains undetectable when uploaded to an open platform and can be selected for model merging, we perform t-SNE analysis [36, 2] on both benign and malicious models. t-SNE excels at preserving the original data distributions in a lower-dimensional space, making it a superior choice for identifying and defending against malicious activity according to previous research [50, 35, 27].
We analyze a total of 80 models—60 benign models fine-tuned on various tasks and 20 malicious models created using LoBAM, also derived from diverse datasets. We apply t-SNE to reduce the dimensionality of model parameters from each layer to three dimensions for visualization purposes. The visualization results for all the layers are depicted in Figure 4 in the Appendix, demonstrating that the parameters of benign and malicious models are indistinguishable within the latent space. This finding provides compelling evidence of the robust concealment capabilities of our proposed attack, confirming its evasion from safety detection.
Defense:  After the model merging process, the model merging creator can employ various defense methods on the merged model to mitigate the attack [49]. To assess whether our attack remains effective in this context, we conduct experiments using well-known defense methods such as Neural Cleanse [37] and Scale-up [11]. The results presented in Table 13 and Table 14 demonstrate that even when defense methods are employed, our attack continues to exhibit strong effectiveness.
Table 12: Attack success rate (%) on on-task and off-task scenarios on dataset combination A, B, and C when all users use ViT-B/16 in the model merging system. SA, TA, Ties, and AM are four different algorithms.
A B C
on-task off-task on-task off-task on-task off-task
SA 100 99.01 99.82 99.24 99.30 99.61
TA 99.35 98.11 100 99.50 98.83 100
Ties 99.59 99.34 99.04 100 99.85 99.13
AM 99.22 99.74 99.52 99.75 97.59 99.18
@@ 6 Conclusion
In this paper, we discovered that existing backdoor attacks on model merging become ineffective due to attackers’ limited computational resources and the resulting reliance on LoRA for fine-tuning pre-trained models. Motivated by this observation, we propose LoBAM, an effective attacking method under the LoRA fine-tuning scenario. LoBAM strategically combines the weights of a malicious and a benign model—each LoRA fine-tuned by the attacker—to amplify attack-relevant components, enhancing the model’s malicious efficacy when deployed in model merging. Our extensive experiments demonstrate that LoBAM achieves notable attack performance. Additionally, our method exhibits excellent stealthiness, making it difficult to detect and defend against using conventional methods. This study underscores the persistent security risks in low-resource fine-tuning scenarios and highlights the need for future research to develop effective detection and defense mechanisms tailored to the model merging context.
Table 13: The result of the Neural Cleanse defense on on-task and off-task scenarios on dataset combination A, B, and C. The attack is successful against the defense if the score is less than 2.
On-task Off-task
A 1.33 1.09
B 1.57 0.83
C 1.04 1.46
Table 14: The result of the Scale-up defense. The attack is considered successful against the defense if the score is less than 0.9.
On-task Off-task
A 0.56 0.39
B 0.32 0.64
C 0.48 0.37